# Learning-based Axial Motion Magnification

### v1

@@ 5 Conclusion
In this work, we present a novel concept, axial motion magnification, which improves the legibility of the motions by disentangling and magnifying the motion representations along axes specified by users. Axial motion magnification can offer easily decipherable motion information in various applications, where recognizing movements along particular axes is crucial. To this end, we propose an innovative learning-based approach for axial motion magnification, incorporating the Motion Separation Module (MSM) to effectively extract and magnify motion representations along two orthogonal orientations. To support this, we establish a new synthetic dataset generation pipeline tailored for axial motion magnification. Our proposed method provides user controllability and significantly enhances the legibility of the motions along chosen axes, showing favorable performance compared to competing methods, even in cases of generic motion magnification.
@@
In this supplementary material, we present the implementation details and additional experiments. We conclude our paper with a discussion section.
@@ Appendix A Implementation Details
We provide the details of data generation pipeline (Sec. A.1) and the loss function for the learning-based axial motion magnification (Sec. A.2).
@@ A.1 Data generation
We randomly sample foreground textures ranging from $7$ to $14$ with segmentation masks from PASCAL VOC [10] and one background from COCO [18]. For each layer, we sample the axial magnification factor $\boldsymbol{\alpha}=(\alpha^{\phi};\alpha^{\phi_{\perp}})$ from the uniform distributions whose values are ranging from $1$ to $80$. Each element of translation parameter $\textbf{d}\in\mathbb{R}^{2}$ is uniformly sampled from the range $-u$ to $u$, where $u$ equals $\mbox{min}(10,30/\mbox{max}(\alpha^{\phi},\alpha^{\phi_{\perp}}))$. It constrains the input motion to be less than $10$ pixels or amplified motions under $30$ pixels. We sample the angle $\phi$ within the range of 00 to $90$ degrees. Note that our method enables axial motion magnification not only in the angle $\phi$ but also in the angle ϕsubscriptitalic-$\phi_{\perp}$, making axial motion magnification possible in 00 to $180$ degrees. To address the loss of subpixel motion due to image quantization, as proposed in DMM [23], we apply uniform quantization noise to the images before quantizing them.
@@
Figure 11: Physical accuracy of our method. We compare the physically calculated sinusoidal wave of pixel displacement (red line) to the $y$-t graph’s waves of 10×10$\times$10 × magnified videos from motion magnification methods. We also provide the $y$-t graph’s wave of the original video and sinusoidal wave before amplification. The $y$-t graph’s wave of Ours matches the actual pixel displacement. The phase-based method [37] exhibits results consistent with the red wave of pixel displacement, albeit suffering from ringing artifacts. Other learning-based methods, such as DMM [23], STM-VMM [16], and Singh et al. [29], also demonstrate correspondences, with a marginal difference in amplification.
@@ A.2 Loss Function
DMM [23] propose the texture loss $L_{\text{texture}}$ and shape loss $L_{\text{shape}}$ to represent intensity and motion information, respectively. These losses are combined with the reconstruction loss $L_{\text{recon}}$, resulting in the loss of DMM $L_{\text{DMM}}$. $L_{\text{DMM}}$ in the case of axial motion magnification is as follows:
$L_{\text{DMM}}$ $=L_{\text{recon}}(\hat{{\mathbf{I}}}^{\phi},\tilde{{\mathbf{I}}}^{\phi})+L_{\text{texture}}(\textbf{T}_{1},\textbf{T}_{2})$
$+L_{\text{shape}}(\textbf{S}^{x}_{2},\acute{\textbf{S}}^{x}_{2})+L_{\text{shape}}(\textbf{S}^{y}_{2},\acute{\textbf{S}}^{y}_{2}).$ (8)
Based on the $L_{\text{DMM}}$, we also integrate perceptual loss $L_{\text{per}}$ [15] and Laplacian loss $L_{\text{lap}}$ into our total loss $L_{\text{total}}$, which corresponds to:
$L_{\text{total}}$ $=L_{\text{recon}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})+\gamma L_{\text{lap}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})+\beta\Bigl{(}L_{\text{texture}}(\textbf{T}_{1},\textbf{T}_{2})$
${\,+}L_{\text{shape}}(\textbf{S}^{x}_{2},\acute{\textbf{S}}^{x}_{2}){\,+}L_{\text{shape}}(\textbf{S}^{y}_{2},\acute{\textbf{S}}^{y}_{2}){\,+}L_{\text{per}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})\Bigr{)},$ (9)
where we set $\beta$ and $\gamma$ to 0.1 and 0.5, respectively. We train our model using two NVIDIA Titan RTX GPUs. The efficacy of $L_{\text{per}}$ and $L_{\text{lap}}$ will be discussed in the Sec. B.5.
@@ Appendix B Additional Experiments
In this section, we assess the physical accuracy (Sec. B.1), motion separation effect of the MSM (Sec. B.2), and the behavior of our method across varying degrees (Sec. B.3). We also demonstrate the compatibility of our method with the temporal filter (Sec. B.4) and conduct ablation studies on the loss function (Sec. B.5).
B.1 Physical Accuracy
To assess the physical accuracy, we examine whether the vibrations of the magnified video, using the proposed method, matched those of actual vibrations. First, we generate a $20$Hz sinusoidal vibration using a vibration generator. Next, we obtain the peak amplitude of acceleration ($\text{m}/\text{s}^{2}$) from the attached accelerometer and convert it into a sinusoidal wave of displacement (m), which is transformed into a sinusoidal wave of pixel displacement (px) on the image plane through pinhole camera geometry. We investigate if this wave corresponds to the vibration of the 10×10$\times$10 × magnified video using the static mode. The transformation from the peak amplitude of acceleration $a$ to the peak amplitude of displacement $\mu$ is as follows:
@@ B.1 Physical Accuracy
where the $\omega$ denotes the frequency of sinusoidal vibration. Using $\mu$, we obtain the sinusoidal wave of real-world displacement $s(t)$ over time $t$ and transform it into pixel displacement $k(t)$, which corresponds to
@@
As shown in Fig. 11, the $y$-t graph’s wave of our method demonstrates a correspondence with the red wave of pixel displacement that is 10×10$\times$10 × amplified. The phase-based method [37] and other learning-based methods [23, 29, 16] also show correspondences, albeit with marginal differences in amplification. These results validate the physical accuracy of our method, as well as that of other motion amplification methods. We list the hyperparameters for converting the acceleration ($\text{m}/\text{s}^{2}$) to pixel displacement (px) in Table 1.
B.2 Motion Separation Effect of the MSM
We assess the effectiveness of the Motion Separation Module (MSM) in distinguishing between two orthogonal directional motions. To explore this, we rotate the video of a vibrator oscillating solely along the $y$-axis by the angle $\phi$ and apply the 10×10$\times$10 × axial motion magnification to the video along the ϕsubscriptitalic-$\phi_{\perp}$ direction using both our method and the modified DMM with the static mode. Subsequently, we compare the time slices in the direction of ϕsubscriptitalic-$\phi_{\perp}$, i.e., the direction with no motion. In this experiment, we set $\phi$ to $30$ degrees. Figure 12 demonstrates the results. Unlike the ϕsubscriptitalic-$\phi_{\perp}$-t graph of the original, where there is no motion in the ϕsubscriptitalic-$\phi_{\perp}$ direction, modified DMM fails to separate the motion and exhibits motion in the ϕsubscriptitalic-$\phi_{\perp}$ direction. In contrast, Ours with MSM effectively separates the motions in two orthogonal directions, showing results similar to the original in the ϕsubscriptitalic-$\phi_{\perp}$ direction.
@@ B.2 Motion Separation Effect of the MSM
Figure 12: Motion separation experiment with MSM. Along the ϕsubscriptitalic-$\phi_{\perp}$ direction, we apply the 10×10$\times$10 × axial motion magnification to the video of a vibrator oscillating only in the $\phi$ direction using Ours and the modified DMM. Unlike the ϕsubscriptitalic-$\phi_{\perp}$-t graph of the original, the modified DMM exhibits vibration in the ϕsubscriptitalic-$\phi_{\perp}$ direction due to the unsuccessful motion separation. In contrast, our method, which uses MSM, effectively separates two orthogonal motions and presents a ϕsubscriptitalic-$\phi_{\perp}$-t graph similar to that of the original.
@@
B.3 Angular Analysis of Axial Motion Magnification
@@
Our learning-based axial motion magnification can magnify the motion along the user-defined direction. We examine whether the behavior of our learning-based axial motion magnification remains consistent with changing angles. As shown in Fig. 13, we rotate the vibrator video at various angles $\phi$ and apply 10×10$\times$10 × axial motion magnification to amplify only the motion corresponding to $\phi$. Subsequently, time slices are obtained from the rotated videos for the same point, and these slices are connected over time. The connected time slices exhibit a smooth transition at boundaries where the angle $\phi$ changes. This demonstrates the consistent behavior of our learning-based axial motion magnification across various angles.
@@
Figure 14: Axial motion magnification with temporal filter. With the Butterworth filter, we 40×40$\times$40 × magnify the rotor imbalance sequence along the $x$-axis, i.e., the axial direction, using Ours and phase-based method [37]. We also show the 40×40$\times$40 × magnified result of DMM [23] with the temporal filter as one reference result of generic motion magnification methods. Our shows the clear magnified output frame and $x$-t graph, which represents legible axial vibrations. The magnified frame of the phase-based method suffers from ringing artifacts, and the DMM demonstrates the magnified frames with artifacts and unclear axial vibrations due to dominant rotational motion, even with the temporal filter.
B.4 Compatibility with the Temporal Filter
Our learning-based axial motion magnification is compatible with temporal filters. To demonstrate compatibility with a temporal filter during axial motion magnification, we apply a 40×40$\times$40 × magnification to the rotor imbalance sequence along the $axial$ direction ($x$-axis) using Ours and phase-based method [37]. We use a Butterworth filter with a frequency range of $3\sim 20$Hz as the temporal filter. The result obtained from DMM [23] with the same filter is provided for comparison with generic motion magnification. As shown in Fig. 14, the magnified frame using our method is free from artifacts, and the $x$-t graph clearly depicts axial vibrations. In contrast, the output of the phase-based method exhibits ringing artifacts, and DMM produces magnified frames with artifacts and unclear axial vibrations due to dominant rotational motion, even with the temporal filter.
In the case of generic motion magnification, our method is also compatible with the temporal filter, as illustrated in Fig. 15. When magnifying the baby sequence 20 times with the IIR filter with a range of $0.04\sim 0.4$Hz, Ours and DMM preserve the boundaries of the baby’s clothing and clearly represent the breathing motion of the baby. The phase-based approach also represents the breathing motion well, but it exhibits slight ringing artifacts.
@@ B.4 Compatibility with the Temporal Filter
Figure 15: Generic motion magnification with temporal filter. We applied generic motion magnification to the baby sequence using the phase-based, DMM, and our methods. The motion magnification factor is set to $20$, and we use the IIR filter as the temporal filter. Ours and DMM preserve the boundaries of the baby’s cloth and depict the breathing motion well. The phase-based method exhibits slight ringing artifacts.
B.5 Ablation Studies of Loss Function
We analyze the effects of $L_{\text{per}}$ and $L_{\text{lap}}$ quantitatively and qualitatively. We progressively introduce $L_{\text{per}}$ and $L_{\text{lap}}$ to $L_{\text{DMM}}$ and evaluate the performance of the networks trained with these additional losses.
Quantitative Analysis
To quantitatively assess the effects of additional losses, we evaluate the performance on both generic and axial evaluation datasets. Figure 16 demonstrates the results. Applying both $L_{\text{per}}$ and $L_{\text{lap}}$ shows superior performance in subpixel and noise tests on both axial and generic evaluation datasets. Adding $L_{\text{per}}$ achieves marginal performance improvement in the two subpixel tests, but the robustness against noises in the generic evaluation dataset decreases. This result is consistent with the observation of DMM [23] that perceptual loss do not yield a noticeable difference. Similarly, using $L_{\text{lap}}$ alone exhibits the trade-offs similar to $L_{\text{per}}$.
Qualitative analysis
We observe that applying both $L_{\text{per}}$ and $L_{\text{lap}}$ is effective to represent the motion of rigid objects. We apply the 40×40$\times$40 × axial motion magnification along the $y$-axis to the rotor imbalance sequence. As shown in Fig. 17, applying both $L_{\text{per}}$ and $L_{\text{lap}}$ preserves the boundary of the shaft well. However, in other cases, the boundary of the shaft is distorted.
Figure 16: Quantitative results on different combinations of losses. Applying both $L_{\text{per}}$ and $L_{\text{lap}}$ shows superior performance in subpixel and noise tests on both axial and generic evaluation datasets. In contrast, applying either $L_{\text{per}}$ or $L_{\text{lap}}$ alone demonstrates a trade-off compared to applying none.
Figure 17: Qualitative results on different combinations of losses. When the rotor imbalance sequence is axially magnified $40$ times along the $y$-axis, applying both $L_{\text{per}}$ and $L_{\text{lap}}$ effectively preserves the boundary of the shaft.
Appendix C Discussion
To improve the legibility of magnified motion, we propose a novel concept, axial motion magnification. Additionally, we propose a new learning-based axial motion magnification method, incorporating a training dataset. Although axial motion magnification serves as one branch that enhances user convenience, another branch can be the method to perform motion magnification in real-time, which is useful and beneficial for various applications of motion magnification. Our method operates at 15.6 FPS on 720p videos, and DMM operates at 18.1 FPS, which does not achieve real-time performance. Improving the efficiency of learning-based axial motion magnification requires further research.

### v3

@@ 5 Conclusion
In this work, we present a novel concept, axial motion magnification, which improves the legibility of the motions by disentangling and magnifying the motion representations along axes specified by users. To this end, we propose an innovative learning-based approach for both axial and generic motion magnification, incorporating the Motion Separation Module (MSM) to effectively extract and magnify motion representations along two orthogonal orientations. To support this, we establish a new synthetic data generation pipeline tailored for axial motion magnification. Our proposed method provides user controllability and significantly enhances the legibility of the motions along chosen axes, showing favorable performance compared to competing methods, even in cases of generic motion magnification. Although axial motion magnification serves as one branch that enhances user convenience, another branch can be the method to perform motion magnification in real-time, which is useful and beneficial for various applications. Similarly to DMM, our method falls short of real-time performance for 720p videos, presenting an avenue for future research in this area.
@@
In this supplementary material, we present the implementation details and additional experiments. Furthermore, we provide axial and generic motion magnification results across various scenarios.
@@ Appendix A Implementation Details
We provide the details of data generation pipeline (Sec. A.1), the loss function for the learning-based axial motion magnification (Sec. A.2) and the details of projection layer (Sec. A.3).
@@ A.1 Data generation
We randomly sample foreground textures ranging from $7$ to $14$ with segmentation masks from PASCAL VOC [10] and one background from COCO [19]. For each layer, we sample the axial magnification factor $\boldsymbol{\alpha}=(\alpha^{\phi};\alpha^{\phi_{\perp}})$ from the uniform distribution whose values are ranging from $1$ to $80$. Each element of translation parameter $\textbf{d}\in\mathbb{R}^{2}$ is uniformly sampled from the range $-u$ to $u$, where $u=\mbox{min}(10,30/\mbox{max}(\alpha^{\phi},\alpha^{\phi_{\perp}}))$. It limits input motions to a maximum of $10$ pixels or ensures amplified motions are kept under $30$ pixels. We sample the angle $\phi$ within the range of 00 to $90$ degrees. Note that our method enables axial motion magnification not only in the angle $\phi$ but also in the angle ϕsubscriptitalic-$\phi_{\perp}$, thus facilitating axial motion magnification within the range of 00 to $180$ degrees. To address the loss of subpixel motion due to image quantization, as proposed in DMM [24], we apply uniform quantization noise to the images before quantizing them.
@@ A.2 Loss Function
DMM [24] proposes the texture loss $L_{\text{texture}}$ and shape loss $L_{\text{shape}}$ to represent intensity and motion information, respectively. These losses are combined with the reconstruction loss $L_{\text{recon}}$, forming the composite loss function of DMM. We slightly modify the loss of DMM to separately impose the loss to the x-axis and y-axis shape representations. The total loss function $L_{\text{total}}$ is as follows:
$L_{\text{total}}$ $=L_{\text{recon}}(\hat{{\mathbf{I}}}^{\phi},\tilde{{\mathbf{I}}}^{\phi})+\beta(L_{\text{texture}}(\textbf{T}_{1},\textbf{T}_{2})+L_{\text{shape}}(\textbf{S}^{x}_{2},\acute{\textbf{S}}^{x}_{2}))+L_{\text{shape}}(\textbf{S}^{y}_{2},\acute{\textbf{S}}^{y}_{2}),$ (9)
where we set $\beta$ to 0.5. We train our model using two NVIDIA Titan RTX GPUs.
A.3 Projection Layer
Figure 11: Projection layer. The projection and inverse projection layers facilitate the synthesis of arbitrarily rotated representations through a linear combination. In (a), the representations aligned with the $x$ and $y$-axes undergo projection onto the $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$ directions. Subsequently, in (b), these representations manipulated within the $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$ directions before being projected back onto the $x$ and $y$-axes.
Motivated by the concept of steerable filters [13], we design the projection layer $P^{\phi}$ and inverse projection layer $P^{-\phi}$ using linear matrices. This enables the synthesis of arbitrarily rotated representations through a linear combination of directional representations. As shown in Fig. 11-(a), the axial shape representation along the canonical $x$ and $y$-axes, which is induced by weight-shared 1D convolutions, are fed to the projection layer $P^{\phi}$. With the linear operation, $P^{\phi}$ projects them and results in the axial shape representations of $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$ directions. Conversely, the inverse projection layer $P^{-\phi}$ projects the outputs of the Manipulator $\Delta^{\phi},\Delta^{\phi_{\perp}}$ back to the canonical $x$ and $y$-axes (Fig. 11-(b)).
@@ Appendix B Additional Experiments
In this section, we assess the physical accuracy on generic motion magnification (Sec. B.1), the physical accuracy of the proposed axial motion magnification (Sec. B.2), motion separation effect of the MSM (Sec. B.3), and the behavior of our method across varying degrees (Sec. B.4). We also demonstrate the per-pixel motion magnification capability of our method (Sec. B.5).
Figure 12: Physical accuracy on generic motion magnification. We compare the physically calculated sinusoidal wave of pixel displacement (red line) to the $y$-t slice’s waves of 10×10$\times$10 × magnified videos from motion magnification methods. We also provide the $y$-t slice’s wave of the original video and sinusoidal wave before amplification for reference. The $y$-t slice’s wave of Ours matches the actual pixel displacement. The phase-based method [39] exhibits results consistent with the red wave of pixel displacement, albeit suffering from ringing artifacts. Other learning-based methods, such as DMM [24], STM-VMM [17], Pan et al[26], and Singh et al. [31], also demonstrate correspondences, with a marginal difference in amplification.
B.1 Physical Accuracy on Generic Motion Magnification
To assess the physical accuracy of each method on generic motion magnification scenario, we examine whether the vibrations of the video which are magnified by each method match those of actual vibrations. First, we generate a $20$Hz sinusoidal vibration using a vibration generator. Next, we obtain the peak amplitude of acceleration ($\text{m}/\text{s}^{2}$) from the attached accelerometer and convert it into a sinusoidal wave of displacement (m), which is transformed into a sinusoidal wave of pixel displacement (px) on the image plane through pinhole camera geometry. We investigate whether this wave corresponds to the vibration of the 10×10$\times$10 × magnified video using the static mode. The transformation from the peak amplitude of acceleration $a$ to the peak amplitude of displacement $\mu$ is as follows:
@@ B.1 Physical Accuracy on Generic Motion Magnification
where $\omega$ denotes the frequency of sinusoidal vibration. Using $\mu$, we obtain the sinusoidal wave of real-world displacement $s(t)$ over time $t$ and transform it into pixel displacement $k(t)$, which corresponds to
@@
As shown in Fig. 12, the sinusoidal wave of our method demonstrates a correspondence with the red wave of pixel displacement that is 10×10$\times$10 × amplified. The phase-based method [39] and other learning-based methods [24, 31, 17, 26] also exhibit correspondences, albeit with slight differences in amplification. These results validate the physical accuracy of our method, as well as that of other motion amplification methods. We provide the hyperparameters for converting acceleration ($\text{m}/\text{s}^{2}$) to pixel displacement (px) in Table 1.
Figure 13: Physical accuracy of the proposed axial motion magnification. (a) Using the Kanade-Lucas-Tomasi (KLT) Tracker [21], we obtain the displacement values of the original video and the video which is 20$\times$ amplified along $x$-axis by our method. (b) We multiply the $x$-axis’ displacement value of the original trajectory by $20$ and compare it with the $x$-axis’ displacement value of the video which is amplified along x-axis by our method. (c) In the $y$-axis direction, we multiply the $y$-axis’ displacement value of the original trajectory by $20$ and compare it with the $y$-axis’ displacement value of the video which is amplified along $x$-axis by our method.
B.2 Physical Accuracy of Axial Motion Magnification
We assess the physical accuracy of axial motion magnification when amplifying the motions, which move in various directions, into only the user-defined direction. As shown in Fig. 13-(a), utilizing the Kanade-Lucas-Tomasi (KLT) Tracker [21], we obtain the displacements of the original video and the video obtained from our method which is magnified $20$ times along the $x$-axis. We evaluate both the physical accuracy and efficacy of axial motion magnification by comparing the displacement values from the trajectory of the video amplified 20×20$\times$20 × using our method against the displacement values obtained by multiplying the original video’s displacement values by $20$. Figure. 13-(b) demonstrates the alignment between the trajectories of the video obtained by our method and the amplified original trajectory. For the y-axis displacement, the direction our method does not aim to amplify, the trajectory of the video obtained by our method aligns with the amplified original trajectory. (Fig. 13-(c)). These observations demonstrate that the proposed axial motion magnification not only preserves physical accuracy but also selectively amplifies motion along user-defined directions.
@@ B.2 Physical Accuracy of Axial Motion Magnification
Figure 14: Motion separation experiment with MSM. Along the ϕsubscriptitalic-$\phi_{\perp}$ direction, we apply the 10×10$\times$10 × axial motion magnification to the video of a vibrator oscillating only in the $\phi$ direction, using both our method and the modified DMM. Contrary to the ϕsubscriptitalic-$\phi_{\perp}$-t slice of the original, the modified DMM exhibits vibration in the ϕsubscriptitalic-$\phi_{\perp}$ direction due to the unsuccessful motion separation. In comparison, our method, leveraging the proposed Motion Separation Module (MSM), successfully distinguishes between the two orthogonal motions, resulting in a ϕsubscriptitalic-$\phi_{\perp}$-t slice that closely resembles the original’s and desired motion trajectory, demonstrating the effectiveness of the MSM.
B.3 Motion Separation Effect of the MSM
We assess the effectiveness of the Motion Separation Module (MSM) in distinguishing between two orthogonal directional motions. To explore this, we rotate the video, where a vibrator oscillates solely along the $y$-axis, by the angle $\phi$ and apply the 10×10$\times$10 × axial motion magnification to the video along the ϕsubscriptitalic-$\phi_{\perp}$ direction using both our method and the modified DMM with the static mode. Subsequently, we compare the time slices in the direction of ϕsubscriptitalic-$\phi_{\perp}$, i.e., the direction with no motion. In this experiment, we set $\phi$ to $30$ degrees. Figure 14 demonstrates the results. Unlike the ϕsubscriptitalic-$\phi_{\perp}$-t slice of the original, where there is no motion in the ϕsubscriptitalic-$\phi_{\perp}$ direction, modified DMM fails to separate the motion and exhibits motion in the ϕsubscriptitalic-$\phi_{\perp}$ direction. In contrast, Ours with MSM effectively separates the motions in two orthogonal directions, showing results similar to the original in the ϕsubscriptitalic-$\phi_{\perp}$ direction.
@@ B.3 Motion Separation Effect of the MSM
B.4 Angular Analysis of Axial Motion Magnification
@@
Our learning-based axial motion magnification can magnify the motion along the user-defined direction. We examine whether the behavior of our learning-based axial motion magnification remains consistent with changing angles. As shown in Fig. 15, we rotate the vibrator video at various angles $\phi$ and apply 10×10$\times$10 × axial motion magnification to amplify only the motion corresponding to $\phi$. Time slices are obtained from the lines that indicate identical positions across the various angle-adjusted videos. Then, the slices are sequentially connected over time. The connected time slices exhibit a smooth transition at boundaries where the angle $\phi$ changes. This demonstrates the consistent behavior of our learning-based axial motion magnification across various angles.
Figure 16: Qualitative results of targeted motion magnification. Our method is capable of per-pixel motion magnification because of the new proposed training dataset. To show this, Given a mask of the notebook, we selectively magnify the motion of the notebook using Ours and Pan et al[26] When targeting the notebook for magnification, we observe that the motion of the air conditioner remains unchanged from the original, while only the motion of the notebook is amplified in Pan et al. and Ours.
B.5 Per-pixel Motion Magnification
During inference, our model demonstrates the ability to perform per-pixel motion magnification, which enables to vary magnification factors across different areas within an image. This capability is endowed by two main components: the angle $\phi$ and object-wise magnification map $\Lambda$, which are main parts of our newly proposed training dataset. We show this spatially selective motion magnification capability by presenting targeted results similar to those achieved by Pan et al[26], which magnify specific objects within an image. Figure 16 displays the targeted motion magnification results of our method, alongside the targeted results obtained by Pan et al. When focusing on magnifying the motion of a notebook, we observe that the motion of the air conditioner remains unchanged from the original footage, while only the motion of the notebook is magnified in both Pan et al. and our method.
Appendix C Additional Results on Diverse Scenarios
In this section, to demonstrate the efficacy of our approach, we present results from diverse scenarios. Our method is capable of both generic and axial motion magnification. Additionally, we observe that the learned shape representations are compatible with the temporal filter, similar to DMM [24]. Therefore, our proposed method provides four configurations based on the motion magnification approach and the application of temporal filters. The following figures demonstrate results on four distinct configurations: axial motion magnification without a temporal filter (Fig. 17), generic motion magnification without a temporal filter (Fig. 18), axial motion magnification with a temporal filter (Fig. 19), and generic motion magnification with a temporal filter (Fig. 20).
Figure 17: Qualitative results of axial magnification. (a) Original: non-magnified. (b) DMM: magnified results with generic method [24]. We magnify $x$-axis motions in air conditioner and $y$-axis motions in gun, and water with (c) phase-based and (d) our methods respectively, plotting $x$-$t$ and $y$-$t$ slices for each of two different points. In cyan scenarios, where magnification aligns with the slice’s axis, ours presents less artifacts and clearer axial vibrations than phase-based, which suffers from severe artifacts and unclear vibrations. In magenta scenarios, when magnification is orthogonal to the slice’s axis, our method isolates motion effectively, preserving time slices similar to (a) without undesired magnification or artifacts. Conversely, DMM and phase-based struggle, leading to time slices deviating from the original, with notable artifacts.
Figure 18: Qualitative results of generic motion magnification. We compare our method to the phase-based [39] method, Singh et al[31], STB-VMM [17], Pan et al[26], and DMM [24] in general motion magnification across various scenarios. Our method demonstrates clear magnified frames and the $x$-t slices.
@@ Appendix C Additional Results on Diverse Scenarios
Figure 19: Axial motion magnification with temporal filter. With the temporal filters, we magnify the rotor imbalance and air conditioner sequences along the $x$-axis, i.e., the axial direction, using (d) Ours and (c) phase-based method [39]. We also show the result of (b) DMM [24] with the temporal filter as one reference result of generic motion magnification methods. In cyan scenarios, where magnification aligns with the slice’s axis, ours shows less artifacts and legible axial vibrations. On the other hand, DMM and phase-based method both suffer from severe artifacts. In addition, DMM shows unclear vibration in the $x$-t slice, even with the temporal filter. In magenta scenarios, when magnification is orthogonal to the slice’s axis, our method effectively isolates the motions which are not aligned with the magnified direction, preserving time slices similar to (a) Original without undesired magnification or artifacts. Conversely, DMM in rotor imbalance sequence and phase-based in air conditioner sequence struggle to disentangle the unwanted motions, which leads to time slices deviating from the original and the magnified frames with artifacts and unclear axial vibrations.
@@
Figure 20: Generic motion magnification with temporal filter. With temporal filters, we applied generic motion magnification to the baby, drum, air conditioner and wheel sequence using the phase-based, DMM [24], Jerk-aware [33] and our methods. Ours and DMM preserve the boundaries of the moving objects while depicting the motion well. The phase-based method exhibits slight ringing artifacts, and the Jerk-aware method shows the unstable separation of the motion signals.